A producer of pens has fixed costs of $17,000 per month which are allocated to the operation and variable costs are $1.80 per pen.

(a) Find the break-even quantity if pens sell at $2.3 each.

(b) Find the profit if the company produces 35,000 pens and pens sell at $2.3 each?
Answer #1

(a) Let the break even quantity be X, then

Total expense = Total revenue

17000 + 1.8*X = 2.3*X

0.5*X = 17000

X = 34000 Units

(b) profit = Revenue - Expense

= 2.3*35000 - (17000 + 1.8*35000)

= $500